# Rejuvenating Cross-Entropy Loss in Knowledge Distillation for Recommender Systems

### v1

@@ 7 Conclusion
Appendix A Relationship between the Rankings Given by the Student and the Teacher
@@ Appendix B Proofs
Then, similar to the proof for Theorem 3 in Bruch et al. (2019), we have
@@ C.1 Experimental Settings
Datasets. We conduct experiments on three public datasets, including CiteULike [^1] https://github.com/changun/CollMetric/tree/master/citeulike-t Wang et al. (2013); Kang et al. (2022; 2021), Gowalla [^2] http://dawenl.github.io/data/gowallapro.zip Cho et al. (2011); Tang & Wang (2018); Lee et al. (2019), and Yelp2018333https://github.com/hexiangnan/sigir16-eals Lee et al. (2019); Kweon et al. (2021).
Following the previous method Xu et al. (2023), we filter out users and items with less than 10 interactions and then split the rest chronologically into training, validation, and test sets in an 8:1:1 ratio. The statistics of the preprocessed datasets are summarized in Table 4.
@@
$\bullet$ CD Lee et al. (2019) samples unobserved items with a ranking-related distribution and uses a point-wise KD loss.
$\bullet$ RRD Kang et al. (2020) adopts a list-wise loss to maximize the likelihood of the teacher’s recommendation list.
$\bullet$ DCD Lee & Kim (2021) corrects what the student has failed to predict with a dual correction loss accurately.
$\bullet$ HetComp Kang et al. (2023) guides the student model by transferring easy-to-hard knwoledge sequences generated from the teacher’s trajectories.
@@ C.3 Ablation Study
This section presents additional results of the ablation study. In Figure 5, we give the results on all KD settings and all datasets. The results suggest similar trends to Figure 3. Specifically, we find that all variants are inferior to the original RCE-KD, demonstrating the effectiveness of all key components. Moreover, RCE-KD w/o S usually performs worse than RCE-KD w/o T. We believe that the reason is that the top items given by the student can exactly satisfy Assumption 4.3, while the top items given by the teacher do not. On the other hand, the superiority of RCE-KD w/ const over RCE-KD w/o T demonstrates the necessity of involving top items from both the student and the teacher. Finally, the superiority of RCE-KD over RCE-KD w/ const and RCE-KD w/o sep validates the effectiveness of our adaptive weighting scheme and the necessity of splitting out the two subsets and treating them separately.
@@ C.4 Hyperparameter Analysis
C.5 Applicability in Sequential Recommendation
@@
Improv.b 5.63% 9.09% 4.08% 3.07% 7.14% 6.10%
p-value 9.6e-5 9.0e-4 3.1e-4 5.5e-4 3.9e-5 1.0e-3
@@
Improv.b 8.99% 4.62% 6.13% 3.38% 5.49% 7.69%
p-value 4.2e-4 9.3e-3 5.3e-4 9.2e-5 7.7e-4 6.5e-3
@@
Our method can be easily applied to other recommendation scenarios, such as sequential recommendation. To verify this, we construct sequential recommendation datasets using the datasets in our paper. Specifically, we take each user’s last interaction as the test item, the second-to-last interaction as the validation item, and the previous interactions as training items.
@@
In Table 8, we report the performance of all methods under four knowledge distillation settings. The results demonstrate that our approach still significantly outperforms all baseline methods on sequence recommendation tasks. Compared to the best baseline method, our method achieves improvements ranging from a minimum of 2.74% to a maximum of 10%. This performance is comparable to our results on top-N recommendation tasks presented in the main text, indicating the strong generalization capability of our method across recommendation tasks.

### v2

@@ 7 Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (No. 62572198 and 92270119).
Appendix A Visualizations
A.1 Relationship between the Rankings Given by the Student and the Teacher
@@
A.2 Tighter NDCG Bound Verification
To empirically verify that our improvement to the CE loss indeed enables us to bound NDCG more tightly, as predicted by our theory, we conduct a comprehensive analysis of the training dynamics across all three recommendation datasets: CiteULike, Gowalla, and Yelp. Specifically, we monitor the NDCG@10 metric on the training set throughout the training process, comparing our improved CE loss with the original CE loss under various teacher-student architecture combinations.
[…]
Higher NDCG values: Across all datasets and teacher-student combinations, RCE-KD consistently achieves higher NDCG@10 values throughout the training process compared to Vanilla CE. This indicates that our approach consistently and stably improves NDCG throughout the entire training process. Moreover, the higher NDCG achieved by our method after training convergence demonstrates its superior ability to optimize NDCG.
Faster growth rate: The training curves reveal that RCE-KD exhibits a steeper learning curve, indicating a more rapid improvement in NDCG during the early training epochs. This accelerated learning can be attributed to the more informative gradient signals provided by RCE-KD. This observation directly supports our theoretical analysis that RCE-KD provides a tighter bound on the ranking metric, as reflected by the superior optimization of NDCG during training.
Faster convergence: In addition to achieving higher peak performance, RCE-KD demonstrates faster convergence to its optimal NDCG value in most cases. The training curves show that RCE-KD reaches its plateau earlier than Vanilla CE, suggesting that the method not only achieves better final performance but also requires fewer training epochs to reach convergence.
Based on the above observations, we can confidently assert that our method indeed achieves tighter bounds on NDCG, as predicted by our theoretical analysis. Consequently, it enables students to learn the ranking capabilities of teachers more effectively, making it more suitable for knowledge distillation in recommender systems.
Figure 5: Training curves of NDCG@10 on the training set for RCE-KD and Vanilla CE across three datasets (CiteULike, Gowalla, Yelp) and five teacher-student architecture combinations.
@@ Appendix B Proofs
Then, similar to the proof for Theorem 3 in (Bruch et al., 2019), we have
@@ C.1 Experimental Settings
Datasets. We conduct experiments on three public datasets, including CiteULike [^1] https://github.com/changun/CollMetric/tree/master/citeulike-t (Wang et al., 2013; Kang et al., 2022; 2021), Gowalla [^2] http://dawenl.github.io/data/gowallapro.zip (Cho et al., 2011; Tang and Wang, 2018; Lee et al., 2019), and Yelp2018333https://github.com/hexiangnan/sigir16-eals (Lee et al., 2019; Kweon et al., 2021).
Following the previous method (Xu et al., 2023), we filter out users and items with less than 10 interactions and then split the rest chronologically into training, validation, and test sets in an 8:1:1 ratio. The statistics of the preprocessed datasets are summarized in Table 4.
@@
$\bullet$ CD (Lee et al., 2019) samples unobserved items with a ranking-related distribution and uses a point-wise KD loss.
$\bullet$ RRD (Kang et al., 2020) adopts a list-wise loss to maximize the likelihood of the teacher’s recommendation list.
$\bullet$ DCD (Lee and Kim, 2021) corrects what the student has failed to predict with a dual correction loss accurately.
$\bullet$ HetComp (Kang et al., 2023) guides the student model by transferring easy-to-hard knwoledge sequences generated from the teacher’s trajectories.
$\bullet$ TARec (Zhuang et al., 2025) proposes a teacher-assisted Wasserstein Knowledge Distillation framework that contains two-stage distillation to bridge the gap between the teacher and student.
@@ C.3 Ablation Study
This section presents additional results of the ablation study. In Figure 6, we give the results on all KD settings and all datasets. The results suggest similar trends to Figure 3. Specifically, we find that all variants are inferior to the original RCE-KD, demonstrating the effectiveness of all key components. Moreover, RCE-KD w/o S usually performs worse than RCE-KD w/o T. We believe that the reason is that the top items given by the student can exactly satisfy Assumption 4.3, while the top items given by the teacher do not. On the other hand, the superiority of RCE-KD w/ const over RCE-KD w/o T demonstrates the necessity of involving top items from both the student and the teacher. The superiority of RCE-KD over RCE-KD w/ const and RCE-KD w/o sep validates the effectiveness of our adaptive weighting scheme and the necessity of splitting out the two subsets and treating them separately. Finally, our method performs even slightly above full-CE in most cases, due to a tighter bound on NDCG than full-item CE (Xu et al., 2024a).
@@ C.4 Hyperparameter Analysis
Figure 9: Training curves of the mean of $s_{u}$ across three datasets (CiteULike, Gowalla, and Yelp).
Figure 10: The histogram of $s_{u}$ at the beginning, middle, and end of training.
C.5 Justification of adaptively scheduling $\gamma$
In this section, we justify our $\gamma$-rule through three principals: (1) decreasing, (2) user-specific, and (3) exponential.
C.5.1 $\mathcal{L}_{1}$’s weight must increase over training
Our distillation objective is
$\mathcal{L}_{\text{RCE-KD}}=(1-\gamma)\mathcal{L}_{1}+\gamma\mathcal{L}_{2}.$ (34)
Let
$G_{1}=\|\nabla\mathcal{L}_{1}\|,\quad G_{2}=\|\nabla\mathcal{L}_{2}\|.$ (35)
Then the gradient contribution of $\mathcal{L}_{1}$ is
$w_{1}(\gamma)=\frac{(1-\gamma)G_{1}}{(1-\gamma)G_{1}+\gamma G_{2}},\quad\frac{dw_{1}}{d\gamma}<0.$ (36)
Thus decreasing $\gamma$ strictly increases the influence of $\mathcal{L}_{1}$.
We compare three schedules: 1) fixed $\gamma$, 2) decreasing $\gamma$, and 3) increasing $\gamma$. The results are given in Table 8.
Table 8: Recommendation performance of different scheduling methods: 1) fixed $\gamma$, 2) decreasing $\gamma$, and 3) increasing $\gamma$.
[…]
Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
[…]
decreasing $\gamma$ 0.0410 0.0174 0.1459 0.1003 0.0634 0.0327
increasing $\gamma$ 0.0372 0.0164 0.1421 0.0990 0.0624 0.0318
[…]
decreasing $\gamma$ 0.0342 0.0160 0.1639 0.1117 0.0674 0.0350
increasing $\gamma$ 0.0335 0.0152 0.1621 0.1106 0.0669 0.0342
[…]
decreasing $\gamma$ 0.0626 0.0341 0.1540 0.1024 0.0726 0.0380
increasing $\gamma$ 0.0603 0.0319 0.1527 0.1008 0.0717 0.0366
We find that only the schedules that increase $\mathcal{L}_{1}$-weight with epoch consistently improve NDCG. This matches our principal of amplifying the influence of $\mathcal{L}_{1}$ and confirms that $\mathcal{L}_{1}$ should progressively dominate as training proceeds.
C.5.2 Why $\gamma$ must be user-specific—and why using $s_{u}$ is justified
In the previous section, $\gamma$ changes with epoch because we analyze global learning dynamics. To enable user-specific adaptivity, however, epoch alone is insufficient: different users progress at different speeds. Thus we define
$s_{u}=\frac{|(Q_{u}^{T})_{1}|}{|Q_{u}^{T}|},$ (37)
which measures how well user $u$ aligns with the teacher.
Figure 9 shows that on average, $s_{u}$ increases monotonically with training proceeding. This implies that making $\mathcal{L}_{1}$-weight increase with $s_{u}$ is consistent with (and indeed equivalent to) the global decision that $\mathcal{L}_{1}$-weight should increase with training proceeding. The crucial difference is that $s_{u}$ captures heterogeneous per-user progress, whereas epoch does not.
Because the distribution of $\{s_{u}\}$ remains highly dispersed during training (as shown in Figure 10), hard users retain low $s_{u}$ while easy users obtain high $s_{u}$. A global $\gamma$ forces both types of users to adopt the same $\mathcal{L}_{1}$$\mathcal{L}_{2}$ weight ratio, which is suboptimal. Thus we adopt the user-specific schedule
$\gamma_{u}=as_{u}+b,\quad\text{L1-weight}=1-\gamma_{u}.$ (38)
Table 9: Performance comparison between global and user-specific scheduling strategies under different teacher–student settings.
[…]
Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
[…]
User-specific 0.0419 0.0181 0.1477 0.1012 0.0648 0.0333
[…]
User-specific 0.0363 0.0164 0.1664 0.1133 0.0699 0.0357
[…]
User-specific 0.0643 0.0350 0.1567 0.1032 0.0734 0.0388
Table 9 shows that this user-specific rule significantly outperforms global $\gamma$.
C.5.3 Why the increase should be exponential
Having established a monotonic increase in $\mathcal{L}_{1}$-weight, we compare two increasing schedules:
  • Linear: $1-\gamma_{\text{lin}}(s_{u})=as_{u}+b$
  • Exponential: $1-\gamma_{\exp}(s_{u})=1-e^{-\beta s_{u}}$
The exponential form provides a phase transition, yielding low weight of $\mathcal{L}_{1}$ initially (allowing $\mathcal{L}_{2}$ to explore) and rapid change later (letting $\mathcal{L}_{1}$ dominate).
Table 10: Performance comparison between linear and exponential scheduling strategies under different teacher–student settings.
[…]
Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
[…]
Exponential 0.0431 0.0194 0.1525 0.1047 0.0667 0.0345
[…]
Exponential 0.0377 0.0171 0.1681 0.1163 0.0716 0.0369
[…]
Exponential 0.0670 0.0366 0.1594 0.1058 0.0754 0.0400
Table 10 shows that the exponential rule consistently outperforms the linear rule, validating its principled nature.
C.5.4 Final Justification
From the analyses above, a proper schedule must:
  1. 1.
    Increase $\mathcal{L}_{1}$-weight over the course of training.
  2. 2.
    Adapt per user, which requires using $s_{u}$.
  3. 3.
    Follow an exponential pattern.
Our final rule is the minimal functional form that satisfies all three principles. Alternative schedules violate at least one of these principles and perform worse.
Appendix D Application on More Recommendation Tasks
Table 11: The Number of transformer blocks (#Block) and number of heads (#Head) for SASRec.
Model CiteULike Gowalla Yelp
#Block #Head #Block #Head #Block #Head
Teacher 4 2 2 4 8 4
Student 1 1 1 1 1 2
@@ Appendix D Application on More Recommendation Tasks
Improv.b 5.63% 9.09% 4.08% 3.07% 5.10% 4.82%
p-value 9.6e-5 9.0e-4 3.1e-4 8.7e-4 7.2e-4 3.8e-4
[…]
Student 0.0054 0.0042 0.0160 0.0086 0.0100 0.0067
CD 0.0073 0.0055 0.0250 0.0169 0.0133 0.0070
RRD 0.0079 0.0057 0.0241 0.0165 0.0129 0.0069
DCD 0.0083 0.0064 0.0269 0.0170 0.0143 0.0074
HetComp 0.0086 0.0066 0.0281 0.0174 0.0136 0.0070
TARec 0.0080 0.0059 0.0271 0.0171 0.0130 0.0067
RCE-KD 0.0092 0.0070 0.0303 0.0181 0.0154 0.0078
Improv.b 6.98% 6.06% 7.83% 4.02% 7.69% 5/41%
p-value 1.3e-3 3.7e-4 6.7e-4 5.5e-5 3.9e-3 9.1e-4
@@
Improv.b 8.99% 4.62% 6.13% 3.38% 5.49% 5.38%
p-value 4.2e-4 9.3e-3 5.3e-4 9.2e-5 6.2e-3 7.4e-4
D.1 Applicability in Sequential Recommendation
@@
Our method can be easily applied to other recommendation scenarios, such as sequential recommendation. To verify this, we construct sequential recommendation datasets using the datasets in our paper. Specifically, we take each user’s last interaction as the test item, the second-to-last interaction as the validation item, and the previous interactions as training items. To further enrich our evaluation, we introduce the classical sequential recommendation model SASRec (Kang and McAuley, 2018) as an additional backbone. We include a new distillation setting in which both the teacher and the student are based on SASRec. The detailed configurations for this setup are provided in Table 11.
@@
In Table 12, we report the performance of all methods under five knowledge distillation settings. The results demonstrate that our approach still significantly outperforms all baseline methods on sequence recommendation tasks. Compared to the best baseline method, our method achieves improvements ranging from a minimum of 2.74% to a maximum of 10%. This performance is comparable to our results on top-N recommendation tasks presented in the main text, indicating the strong generalization capability of our method across recommendation tasks.
D.2 Applicability in Multi-Modal Recommendation
Table 13: Statistics of datasets with multi-modal item Visual (V), Acoustic (A), Textual (T) contents.
Dataset Netflix Tiktok Electronics
Modality V      T V      A      T V       T
Feat. Dim. 512   768 128   128   768 4096   384
User 43,739 14,343 41,691
Item 17,239 8,690 21,479
Interaction 609,341 276,637 359,165
Sparsity 99.919% 99.778% 99.960%
Table 14: Dimensions of models on multi-modal recommendation task.
Dataset Netflix Tiktok Electronics
Model BM3 MF LightGCN BM3 MF LightGCN BM3 MF LightGCN
Dimension 400 20 20 200 20 20 200 20 20
Table 15: Recommendation performance on multi-modal recommendation task. The best results are in boldface, and the best baselines are underlined. Improv.b denotes the relative improvement of RCE-KD over the best baseline. LGCN stands for LightGCN. A paired t-test is performed over 5 independent runs for evaluating $p$-value ($\leq 0.05$ indicates statistical significance).
[…]
Recall@20 NDCG@20 Recall@20 NDCG@20 Recall@20 NDCG@20
[…]
Student 0.1420 0.0498 0.0417 0.0153 0.0182 0.0073
HetComp 0.1783 0.0755 0.0659 0.0238 0.0320 0.0133
PromptMM 0.1732 0.0728 0.0637 0.0221 0.0297 0.0120
TARec 0.1792 0.0760 0.0656 0.0230 0.0333 0.0139
RCE-KD 0.1884 0.0797 0.0702 0.0253 0.0347 0.0146
Improv.b 5.13% 4.87% 6.53% 6.30% 4.20% 5.04%
p-value 2.0e-3 5.9e-4 1.7e-3 9.7e-5 6.6e-3 2.8e-4
[…]
Student 0.1507 0.0537 0.0562 0.0189 0.0211 0.083
HetComp 0.1811 0.0747 0.0629 0.0220 0.0317 0.0129
PromptMM 0.1779 0.0728 0.0603 0.0215 0.0323 0.0137
TARec 0.1834 0.0753 0.0651 0.0235 0.0341 0.0144
RCE-KD 0.1907 0.0804 0.0705 0.0255 0.0359 0.0153
Improv.b 3.98% 6.77% 8.29% 8.51% 5.28% 6.25%
p-value 7.9e-4 4.0e-3 9.4e-4 5.9e-3 7.3e-4 8.2e-5
To further examine whether our approach generalizes to multi-modal recommendation scenarios, we conduct experiments on three widely used multi-modal recommendation datasets:
  • Netflix: The Netflix dataset (Wei et al., 2024) provides user–item interaction logs collected from the streaming platform. Each movie is accompanied by multimodal metadata, most notably poster images and titles. Visual representations are derived using the CLIP-ViT encoder (Radford et al., 2021), while textual information is transformed into embeddings via a pre-trained BERT model (Devlin et al., 2019).
  • Tiktok: The Tiktok micro-video dataset (Wei et al., 2023) offers interaction histories together with three complementary types of content: visual, acoustic, and textual. Visual and audio signals from short videos are processed into 128-dimensional feature vectors, and the accompanying captions are embedded using Sentence-BERT (Reimers and Gurevych, 2019).
  • Electronics: The Amazon Electronics dataset (He and McAuley, 2016; McAuley et al., 2015) consists of user reviews and item metadata from the consumer electronics category. Images are represented using 4,096-dimensional features extracted by pre-trained CNN-based models (He and McAuley, 2016). Textual information—formed by integrating titles, descriptions, category labels, and brand attributes—is encoded as 384-dimensional embeddings using Sentence-BERT (Reimers and Gurevych, 2019).
In this task of knowledge distillation for multi-modal recommender systems, we compare our method with three baselines. Specifically, we select two state-of-the-art knowledge distillation methods for multi-modal recommender systems: PromptMM (Wei et al., 2024) and TARec (Zhuang et al., 2025). PromptMM introduces soft prompt-tuning, together with a disentangled multi-modal list-wise distillation and modality-aware re-weighting mechanism. TARec proposes a teacher-assisted Wasserstein Knowledge Distillation framework that contains two-stage distillation to bridge the gap between the teacher and student. We also utilize HetComp (Kang et al., 2023), the state-of-the-art knowledge distillation method for general recommender systems as a baseline. Note that our method and HetComp rely solely on the teacher’s logits for distillation, whereas PromptMM and TARec use both logits and embeddings. To ensure a fair comparison, we disable their embedding-level distillation components and retain all other parts of their methods.
[…]
Table 15 summarizes the experimental results. In all settings, RCE-KD significantly improves over the students and consistently outperforms all baseline methods. In most cases, our method approaches the teacher’s performance, remaining just slightly lower, or even slightly higher, while other methods are farther behind. These results confirm that our distillation approach remains effective when transferring knowledge from multi-modal teachers to lighter ID-based students. Our experimental results demonstrate that the proposed method is not limited to distillation within ID-based models.